### v1

# Golden Touchstone: A Comprehensive Bilingual Benchmark for Evaluating Financial Large Language Models

Abstract
As large language models become increasingly prevalent in the financial sector, there is a pressing need for a standardized method to comprehensively assess their performance. However, existing finance benchmarks often suffer from limited language and task coverage, as well as challenges such as low-quality datasets and inadequate adaptability for LLM evaluation. To address these limitations, we propose "Golden Touchstone", the first comprehensive bilingual benchmark for financial LLMs, which incorporates representative datasets from both Chinese and English across eight core financial NLP tasks. Developed from extensive open source data collection and industry-specific demands, this benchmark includes a variety of financial tasks aimed at thoroughly assessing models’ language understanding and generation capabilities. Through comparative analysis of major models on the benchmark, such as GPT-4o, Llama3, FinGPT and FinMA, we reveal their strengths and limitations in processing complex financial information. Additionally, we open-sourced Touchstone-GPT, a financial LLM trained through continual pre-training and financial instruction tuning, which demonstrates strong performance on the bilingual benchmark but still has limitations in specific tasks.This research not only provides the financial large language models with a practical evaluation tool but also guides the development and optimization of future research. The source code for Golden Touchstone and model weight of Touchstone-GPT have been made publicly available at https://github.com/IDEA-FinAI/Golden-Touchstone, contributing to the ongoing evolution of FinLLMs and fostering further research in this critical area.
{CJK*}
UTF8gbsn
11footnotetext: Equal Contribution.22footnotetext: Work done during internship at IDEA Research.33footnotetext: Corresponding Author.footnotetext: Correspondence to: Jian Guo <guojian@idea.edu.cn>.
1 Introduction
The rapid development of both proprietary(Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Anthropic, 2024; Team et al., 2023) and open-source Large Language Models (LLMs) has led to their increasing importance and application across various fields(Touvron et al., 2023a, b; AI@Meta, 2024; Bai et al., 2023; Yang et al., 2024a; DeepSeek-AI, 2024; Young et al., 2024; Zeng et al., 2023; Baichuan, 2023; Gan et al., 2023; Zhang et al., 2022) , including finance(Wu et al., 2023; Lopez-Lira and Tang, 2023), healthcare Thirunavukarasu et al. (2023); Tian et al. (2023), and law Cui et al. (2023); Xiao et al. (2021). Among these, the financial sector stands out as one of the most important areas for LLM application, due to its rich textual information and high practical value.
In recent years, a variety of advanced financial large language models (FinLLMs) have emerged, capable of specialized tasks such as financial sentiment analysis, content summarization, stock movement prediction, and specialized question answering, as depicted in Figure.1. These models, including FinGPT (Yang et al., 2023), PIXIU Xie et al. (2023), CFGPT (Li et al., 2023), DISC-FinLLM (Chen et al., 2023), and XuanYuan (Zhang and Yang, 2023), leverage unique frameworks and tuning methods to enhance their performance on domain-specific benchmarks, offering robust solutions for real-world financial applications. Alongside the rapid emergence of FinLLMs, there has also been a significant increase in financial benchmarks. In the English financial domain, FLUE(Shah et al., 2022) was the first publicly available benchmark for assessing English financial language understanding. Similarly, BBT(Lu et al., 2023a) was the pioneering open-source financial NLP benchmark in Chinese. As LLMs and financial applications have continued to develop, more financial benchmarks have been introduced(Xie et al., 2023, 2024; Yang et al., 2023; Lei et al., 2023; Zhang et al., 2023). While these benchmarks have provided valuable resources and evaluation criteria for assessing current LLMs in finance, there remain challenges, such as the presence of low-quality or unsuitable for LLMs datasets within certain tasks, leading to generally poor evaluation results across various models. Furthermore, Current benchmarks also suffer from insufficient language and task coverage, which prevents a comprehensive evaluation of large financial language models.
Figure 1: Financial large language models are designed to perform specialized tasks such as financial sentiment analysis, content analysis, stock movement prediction, and financial analyst level question answering by interpreting and processing structured instructions and various input data to generate precise outputs.
To address these gaps, we propose Golden Touchstone, a bilingual financial benchmark that consolidates representative datasets across eight financial NLP tasks in both Chinese and English. Golden Touchstone provides high-quality datasets, task-aligned metrics, and instructional templates tailored to guide LLMs in generating task-appropriate responses. The Golden Touchstone organizes all task data into an instruction, input, and output format, facilitating the use and evaluation of different models. We evaluated current open-source general-purpose LLMs and FinLLMs using the Golden Touchstone benchmark. Results indicate that while GPT-4o, Qwen-2, Llama-3, and FinMA performed well across several tasks like financial sentiment analysis and entity extration, there remains considerable room for improvement in areas like credit card scoring dataset within classification tasks and stock movement prediction. Additionally, we have open-sourced Touchstone-GPT, a model trained through domain-specific continual pre-training and financial instruction tuning. It achieved acceptable performance on the bilingual benchmark but showed limitations in tasks such as stock movement prediction and question answering. These findings underscore the need for more high-quality training data and potentially more suitable model architectures for finance-specific applications. Our main contributions are as follows:
  • Introduction of Golden Touchstone, the first comprehensive bilingual benchmark for financial Large Language Models (LLMs), encompassing 22 datasets across eight tasks in both Chinese and English. Compared to previous FinLLMs’ benchmarks, our benchmark offers enhanced diversity, systematicity, and adaptability.
  • A thorough evaluation of state-of-the-art LLMs and FinLLMs, including GPT-4o, Qwen-2, Llama-3, FinGPT and FinMA and so on, on the Golden Touchstone. This evaluation highlights the key strengths and shortcomings in model performance across various tasks, and suggests directions for future research in financial LLMs.
  • The open-source release of Touchstone-GPT, a specialized financial LLM that has undergone domain-specific continuous pre-training and instruction tuning. This model serves as a new baseline for subsequent FinLLMs research, fostering further advancements in Financial AI.
2 Related Works
2.1 Financial Large Language Models
In recent years, large language models (LLMs) tailored for the financial domain have become increasingly important. Several significant FinLLMs have been developed: BloombertGPT (Wu et al., 2023) marks the beginning of the FinLLM era. FinGPT (Yang et al., 2023) is an open-source framework designed for financial LLMs. It focuses on a data-centric approach, offering transparent resources for developing FinLLMs. The framework highlights the importance of an automated data curation pipeline and lightweight low-rank adaptation techniques. The follow-up work, FinGPT-Forecaster, continues to use LLM for stock price prediction, which is not ideal for numerical computation tasks. PIXIU (Xie et al., 2023) provides a comprehensive framework that includes the first financial LLM fine-tuned on LLaMA with instruction data, a 136K instruction dataset for fine-tuning, and an evaluation benchmark with 5 tasks and 9 datasets. This framework supports the evaluation and development of financial LLMs. CFGPT (Li et al., 2023) introduces a Chinese Financial Generative Pre-trained Transformer framework. It consists of CFData, a dataset for pre-training and supervised fine-tuning; CFLLM, a financial LLM capable of processing financial texts; and CFAPP, a deployment framework for real-world financial applications. DISC-FinLLM Chen et al. (2023) is based on a Multiple Experts Fine-tuning Framework. It enhances general LLMs by adding multi-turn question answering, domain text processing, mathematical computation, and retrieval-enhanced generation capabilities. XuanYuan 2.0 (Zhang and Yang, 2023) is the largest Chinese chat model to date, built on the BLOOM-176B architecture. It introduces a novel training method called hybrid-tuning to address catastrophic forgetting, combining general-domain and domain-specific knowledge. FinMem (Yu et al., 2023), FinRobot (Yang et al., 2024c), and FinReport (Li et al., 2024) predict stock price trends using agents and tools, such as quantitative models that use multiple alpha factors and alternative factor models that analyze news and textual data. FinVis-GPT (Wang et al., 2023b) and FinTral Bhatia et al. (2024) develop multi-modal financial models, similar to other visual language models (VLM) (Zhu et al., 2023; Liu et al., 2023; Lu et al., 2023b). These models represent the cutting-edge in FinLLMs, each contributing unique approaches and advancements to the field of AI in finance.
2.2 Benchmarks for FinLLMs
In the current landscape of open-source financial evaluation benchmarks, FLUE (Shah et al., 2022) was the first open-source benchmark introduced for financial language understanding in English. It includes five financial tasks: Financial Sentiment Analysis (Malo et al., 2014; Maia et al., 2018b), Financial News Classification (Sinha and Khandait, 2021), Structure Boundary Detection (FinSBD3, 2021), Named Entity Recognition (Alvarado et al., 2015), and Question Answering (Maia et al., 2018a). Building on FLUE, FinGPT (Yang et al., 2023) introduced the Financial Relation Extraction task with the addition of the FinRED dataset (Sharma et al., 2022), while also expanding the sentiment analysis task by incorporating TFNS (Magic, 2022) and NWGI (Yang, 2023) datasets. Additionally, FinGPT created prompt-based instructions across all datasets, culminating in the Financial Instruction Tuning Benchmark. Around the same time, other open-source benchmarks, such as PIXIU (Xie et al., 2023) and FinBen (Xie et al., 2024), were developed. PIXIU includes a wide range of financial tasks and datasets, while FinBen, expanding on PIXIU, introduced numerous supplementary financial datasets to evaluate LLMs comprehensively and gauge their proficiency across various financial scenarios.In the Chinese domain, BBT-Benchmark (Lu et al., 2023a) was the first Chinese financial benchmark, covering multiple financial NLP tasks, including sentiment analysis, news classification, summarization, relation extraction, and question answering. Later, FinEval (Zhang et al., 2023) benchmark emerged, providing an evaluation framework for LLMs in financial contexts, encompassing multiple-choice questions across finance, economics, accounting, and certification. DISC-FinLLM (Chen et al., 2023) evaluated financial LLMs on top of these benchmarks, introducing additional tasks focused on financial calculations and current events to assess LLMs in applied financial scenarios. The most recent Chinese Financial benchmark, CFBenchmark (Lei et al., 2023), introduced three key tasks like Financial Entity Recognition, expanding evaluation resources for Chinese financial tasks.
The increasing number of benchmarks in both English and Chinese financial domains has established valuable evaluation frameworks for LLMs in finance, though challenges remain. For example, data quality across various English financial datasets is inconsistent, limiting LLMs from reaching their full potential, as seen in FinBen and PIXIU, where low-quality datasets yielded suboptimal performance metrics. In Chinese benchmarks, tasks are often skewed toward multiple-choice questions, with limited alignment with English financial tasks. Moreover, no bilingual benchmark currently exists for comprehensive LLM evaluation in financial NLP tasks across English and Chinese. In this work, we integrate representative datasets from both languages, creating FinLLM-benchmark. We compile and align eight financial NLP tasks in English and Chinese, selecting high-quality datasets for each task and appropriate metrics to effectively assess model performance.
3 Benchmark Design
3.1 Current Benchmark Status
In current open-source financial benchmarks, FLUE(Shah et al., 2022) is among the most widely used benchmarks for English financial NLP tasks, encompassing five financial tasks: financial sentiment analysis, news headline classification, named entity recognition, structural boundary detection, and question answering. FinGPT(Yang et al., 2023) extended this benchmark with additional financial tasks. PIXIU(Xie et al., 2023) proposed a more comprehensive benchmark, covering more financial tasks and datasets than FLUE and introducing a stock movement prediction task. Building on PIXIU, FinBen(Xie et al., 2024) added numerous financial datasets to thoroughly evaluate large language models (LLMs) and assess their proficiency across diverse financial scenarios. In the Chinese financial domain, BBT-Benchmark(Lu et al., 2023a) was the first Chinese benchmark for financial evaluation, covering financial news classification, summarization, relation extraction, question answering, negative news detection, and sentiment classification for financial social media texts. Additionally, the FinEval(Zhang et al., 2023) benchmark provides an evaluation framework for LLM knowledge in finance, including multiple-choice datasets across finance, economy, accounting, and certification fields. Subsequently, the FinanceIQ(Zhang and Yang, 2023) and CFBenchmark(Lei et al., 2023) benchmarks were introduced, with FinanceIQ including a broader range of Chinese financial multiple-choice datasets, while CFBenchmark focuses on three major financial task categories: financial entity recognition, financial text classification, and financial content generation.
However, data quality issues in certain tasks within these benchmarks—characterized by inconsistent, unstructured input information—often hinder the accurate comprehension and response generation of financial LLMs. Large language models currently face challenges in numerical understanding and computation, which may be attributed to multiple factors such as tokenizers, positional encoding, and others (Shen et al., 2023; Akhtar et al., 2023; Schwartz et al., 2024). Therefore, using specialized agents is often a more ideal approach for handling these tasks rather than relying on vanilla large language models. This issue is especially evident in PIXIU’s credit analysis, question answering, and stock movement prediction tasks, where model performance metrics are frequently low. Additionally, varying label formats across datasets for the same task make it challenging to match model outputs with labels directly. Notably, most current financial benchmarks are monolingual, lacking bilingual evaluation options to assess financial models comprehensively in both English and Chinese.
Table 1: Diversity of Financial Analysis Tasks Across Different Financial Large Language Model Benchmarks
Benchmarks Sent. Anal. Classif. Ent. Extr. Rel. Extr. Multi. Choice Summ. Quest. Ans. Stock Pred.
FinGPT-Bench (Wang et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
FinBen (Xie et al., 2024) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BBT-Fin (Lu et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Fin-Eval (Zhang et al., 2023) $\checkmark$
FinanceIQ (Zhang and Yang, 2023) $\checkmark$
CFBenchmark (Lei et al., 2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Golden-Touchstone $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 2: Language Coverage, Systematicity, Adaptability, and Model Training Stage for Benchmarks. Systematicity refers to whether benchmarks are established according to a comprehensive system standard. Adaptability indicates whether the tasks are suitable for large language models.
Benchmarks Language Systematicity Adaptability Model Training
EN CN Cont. Pre-train Instr. Tuning
FinGPT-Bench (Wang et al., 2023a) $\checkmark$ Medium High $\checkmark$
FinBen (Xie et al., 2024) $\checkmark$ High Medium $\checkmark$
BBT-Fin (Lu et al., 2023a) $\checkmark$ Medium High $\checkmark$
Fin-Eval (Zhang et al., 2023) $\checkmark$ High High
FinanceIQ (Zhang and Yang, 2023) $\checkmark$ Medium High $\checkmark$ $\checkmark$
CFBenchmark (Lei et al., 2023) $\checkmark$ High High $\checkmark$ $\checkmark$
Golden-Touchstone $\checkmark$ $\checkmark$ High High $\checkmark$ $\checkmark$
To address these limitations, we propose a unified benchmark that incorporates representative financial datasets in both English and Chinese. By selecting high-quality, authoritative datasets and optimizing inputs—such as using news-based input for stock trend prediction rather than tabular data—we aim to make tasks more compatible with LLM assessment. This bilingual benchmark will facilitate model evaluation on both English and Chinese financial tasks, align financial NLP tasks across languages, and provide a holistic measure of financial model performance across linguistic contexts.
The financial sub-tasks included in previous financial benchmarks are shown in Table.1, which illustrates the diversity of financial analysis tasks across different benchmarks. Table.2 provides an overview of the language coverage, systematicity, adaptability, and model training of these benchmarks. In this context, systematicity refers to whether benchmarks are established according to a comprehensive and standardized framework, while adaptability assesses whether the tasks are suitable for large language models. As seen in Table 1, the task coverage of earlier benchmarks is not fully comprehensive, with some financial tasks missing. To address this, we selected representative datasets from various open-source benchmarks based on the task categories in Table 1, ensuring a broad diversity of tasks in the Golden-Touchstone benchmark. Therefore, Golden-Touchstone includes tasks that broadly cover those in previous benchmarks. From Table 2, it is evident that most current financial benchmarks are monolingual, with inconsistent levels of systematicity and adaptability. In response, we have compiled the first bilingual financial benchmark from open-source datasets, achieving high standards in terms of both systematicity and adaptability.
3.2 Golden Touchstone Benchmark Design
Figure 2: Financial NLP tasks are categorized along two dimensions: task types, divided into financial NLU (Natural Language Understanding) and financial NLG (Natural Language Generation), and language, categorized as English and Chinese. We organized the collected high-quality datasets along these axes.
Golden Touchstone benchmark redefines popular open-source financial benchmarks, as illustrated in Figure 2. We classify financial NLP tasks by type—financial NLU (Natural Language Understanding) and financial NLG (Natural Language Generation)—and by language, distinguishing between English and Chinese datasets. These dimensions guide the organization of high-quality open-source datasets. NLU tasks include Financial Sentiment Analysis, Financial Knowledge Examination, Financial Information Extraction, and Financial Text Understanding, while NLG tasks include Financial Summary Analysis and Financial Question Answering. Collectively, these cover eight sub-tasks: Sentiment Analysis, Classification, Entity Recognition, Relation Extraction, Multiple Choice, Summarization, Question Answering, and Stock Movement Prediction. For the English benchmark, tasks were selected primarily from the PIXIU(Xie et al., 2023) and FinBen(Xie et al., 2024) financial benchmarks, covering sentiment analysis, classification, entity recognition, question answering, stock movement prediction, summarization, and credit scoring. Due to low data quality and inconsistent labeling formats in some datasets, we excluded certain lower-quality data while retaining those with relatively higher quality. FinBen’s credit scoring task, a binary classification task, was merged into the general classification task. Additionally, as the stock movement prediction task in FinBen uses time-series tabular data, which tends to be ineffective for prediction in large language models (LLMs) due to LLMs’ limited sensitivity to numerical data in inference contexts, we replaced this task with the DJIA (Aaron7Sun, 2016) dataset, using daily news to predict Dow Jones Index trends. To ensure alignment with the Chinese benchmark tasks, we also incorporated the Financial Relation Extraction task from FinGPT (Yang et al., 2023) and the Multiple Choice task from the CPA-Eval (Yang et al., 2024b) dataset, using the validation set as the test set due to the absence of labels in the test set. For the Chinese benchmark, tasks were chosen primarily from the BBT-benchmark (Lu et al., 2023a) and FinEval benchmark (Zhang et al., 2023) , including sentiment analysis, classification, entity extraction, question answering, summarization, relation extraction, and multiple-choice tasks. As test sets for these benchmarks lack labels, we used the validation set for evaluation. To expand the multiple-choice tasks, we added the CPA (Yang et al., 2024b) dataset, using its validation set as a test set, also without labels. For alignment with the English benchmark tasks, we introduced stock movement prediction with the stockA (Zou et al., 2022) dataset, which includes individual stock news and stock data for five days post-publication. The price movement label is based on the difference between the final day’s closing price and the initial net asset value, enabling LLM-based predictions on stock trends informed by news. Dataset details are provided in Table.3 and Table.4.
Table 3: Overview of English Finance Evaluation Datasets by Task Type, Sample Sizes (Training, Validation, Test), and Evaluation Metrics
Task Dataset Train Valid Test Metrics
Sentiment Analysis FPB 3100 776 970 Weighted-F1 ACC
FiQA-SA 750 188 235 Weighted-F1 ACC
Classification Headlines 71900 10300 20500 Weighted-F1 ACC
FOMC 1984 - 496 Weighted-F1 ACC
lendingclub 9417 1345 2691 Weighted-F1 MCC
Entity Recognition NER 408 103 98 Entity-F1
Relation Extraction FinRE 27558 - 5112 Relation-F1
Multiple Choice CFA 1884 100 20 Weighted-F1 ACC
Summarization EDTSUM 8000 - 2000 ROUGE BLEU
Question Answering FinQa 6251 883 1147 RMACC
ConvfinQa 8890 2210 1490 RMACC
Stock Movement Prediction DJIA 1591 - 398 Weighted-F1 ACC
Table 4: Overview of Chinese Finance Evaluation Datasets by Task Type, Sample Sizes (Training, Validation, Test), and Evaluation Metrics
Task Dataset Train Valid Test Metrics
Sentiment Analysis FinFE-CN 16157 2020 2020 Weighted-F1 ACC
Classification FinNL-CN 7071 884 884 ORMACC
Entity Extraction FinESE-CN 14252 1781 1782 ORMACC
Relation Extraction FinRE-CN 13486 1489 3727 RMACC
Multiple Choice FinEval 1071 170 3340 Weighted-F1 ACC
CPA 6268 1444 6 Weighted-F1 ACC
Summarization FinNA-CN 28800 3600 3600 ROUGE BLEU
Question Answering FinQa-CN 19906 2469 2480 RMACC
FincQa-CN 21965 2741 2745 RMACC
Stock Movement Prediction AStock 11815 1477 1477 Weighted-F1 ACC
English Finance Benchmark: We constructed this benchmark primarily based on the PIXIU and FinBen Benchmark. For sentiment analysis, we selected the FPB (Malo et al., 2014) and FiQA-SA (Maia et al., 2018b) datasets, evaluated using accuracy (ACC) and weighted F1 score (weighted-F1), two widely adopted classification metrics. For classification tasks, we used the Headlines (Sinha and Khandait, 2021), FOMC (Shah et al., 2023), and LendingClub (Feng et al., 2023) datasets. Evaluation of the Headlines and FOMC datasets was conducted with ACC and weighted-F1, while the LendingClub dataset was assessed using Matthews Correlation Coefficient (MCC) and accuracy to align with FinBen standards. For entity extraction, we used the NER (Alvarado et al., 2015) dataset, employing entity-level F1 score (Entity-F1) in line with the PIXIU framework. The EDTSUM (Zhou et al., 2021) dataset was selected for summarization tasks, evaluated using Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002) and ROUGE scores (Lin, 2004), which are standard metrics in text generation. For question answering, we chose the FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) datasets, opting for Regex Match Accuracy (RMACC) over exact match accuracy, as RMACC better accommodates our general-purpose models that may not directly output exact answers. For relation extraction, we used the FinRED (Sharma et al., 2022) dataset, evaluated with the Relation-F1 metric, consistent with FinGPT. The CFA (Yang et al., 2024b) dataset was used for multiple-choice tasks, evaluated using weighted-F1 and ACC. Finally, we selected the DJIA (Aaron7Sun, 2016) dataset for stock movement prediction, evaluated using ACC and weighted-F1. Unlike PIXIU’s use of MCC for this task, we employed weighted-F1 due to generation models’ tendency to produce low-accuracy [unknown] labels.
Chinese Finance Benchmark: We constructed this benchmark primarily based on the BBT-Benchmark (Lu et al., 2023a) and FinEval Benchmark (Zhang et al., 2023). For sentiment analysis, we used the FinFE-CN dataset, evaluated by weighted-F1 and ACC. For question answering, we used the FinQA-CN and FinCQA-CN datasets, assessed using RMACC. Classification tasks employed FinNL-CN, requiring full matches for multi-answer questions and evaluated by Overall Regex Match Accuracy (ORMACC).For relation extraction, we used the FinRE-CN dataset, also evaluated with RMACC. The entity extraction task employed the FinESE-CN dataset, evaluated using ORMACC due to the dataset’s lack of BIO labels for original text entities. Summarization tasks used the FinNA-CN dataset, evaluated with BLEU and ROUGE scores in line with the English benchmark. For multiple-choice tasks, we used FinEval and CPA (Yang et al., 2024b), with FinEval combining finance, economy, accounting, and certification questions, both evaluated with weighted-F1 and ACC. The AStock (Zou et al., 2022) dataset was used for stock movement prediction, assessed by ACC and weighted-F1.
In tasks with [unknown] labels, such as classification, multiple choice, and stock prediction, we treated [unknown] labels as incorrect to maintain consistency. This adjustment proved more reliable than alternative methods like sample omission, random assignment, or fixed selection, which may introduce bias.
4 Experiments
4.1 Experimental Setup
Baselines. We conducted an extensive experimental evaluation against the Golden Touchstone Benchmark, incorporating a comprehensive array of models. For all models and inference tasks, we set the PyTorch and CUDA random seeds and configured the model with a greedy decoding strategy. This ensures reproducibility of experimental results and eliminates the influence of sampling decoding strategies on the final generated outputs. This included cutting-edge commercial models such as GPT-4o (OpenAI, 2023), alongside prominent open-source alternatives like Meta Llama-3 (AI@Meta, 2024) and Alibaba Qwen-2 (Yang et al., 2024a). Additionally, we integrated the latest and most influential financial language models (FInLLMs), namely FinGPT (Yang et al., 2023), FinMA (Xie et al., 2024), CFGPT (Li et al., 2023), and DISC-FinLLM Chen et al. (2023). These models were meticulously selected to represent a diverse spectrum of capabilities, ranging from general-purpose language understanding to specialized financial domain expertise. Our experiments aimed to rigorously assess the performance, robustness, and adaptability of each model within the context of financial data processing and analysis. The results provide valuable insights into the strengths and limitations of current state-of-the-art models, offering a foundation for future advancements in financial language modeling.
Touchstone-GPT Training. To further contribute to the research and development of FInLLMs and Financial benchmarks for LLMs, we have meticulously trained and open-sourced a Touchstone-GPT model. This initiative aims to serve as a valuable resource for advancing the field, providing a robust and versatile model that can be utilized for a wide range of financial language tasks. By making the Touchstone-GPT model publicly available, we aspire to foster collaboration and innovation within the research community, facilitating the development of more sophisticated and effective financial language models. Our hope is that this open-source contribution will not only enhance the current state of research but also inspire new methodologies and benchmarks that push the boundaries of what is possible in financial language modeling. We adopted a two-stage training strategy comprising continuous pre-training and post-training, based on the Qwen-2 (Yang et al., 2024a) foundational model. During the continuous pre-training phase, we initially conducted pre-training on a high-quality financial corpus containing 100 billions tokens, which included textbooks, encyclopedias, research reports, news articles, and real-time analysis, all meticulously cleaned. In the post-training phase, we employed a standard instruction fine-tuning strategy, collecting, cleaning, and formatting a high-quality dataset of 300,000 instruction-response pairs shown in Tabel.3 and Table.4. To avoid catastrophic forgetting in general tasks, we also incorporated general-domain pre-training corpora (Gan et al., 2023) and instruction-tuning corpora (Peng et al., 2023) into continuous pre-training and post-training. This culminated in the final Touchstone-GPT model. We utilized Megatron(Shoeybi et al., 2019) for continuous pre-training and LlamaFactory(Zheng et al., 2024) for instruction post-training as our training frameworks, respectively. In this study, we employ an advanced model training setup using the AdamW optimizer (Kingma and Ba, 2014) with a learning rate of 1.0e-5, cosine annealing scheduler (Szegedy et al., 2016), and a 10% warmup ratio (He et al., 2016; Goyal et al., 2017) to enhance training stability and convergence. We enable gradient accumulation (Shoeybi et al., 2019) and checkpointing (Chen et al., 2016) to simulate larger batch sizes and reduce memory footprint. Training is conducted in mixed bfloat16 precision (Micikevicius et al., 2017; Wang and Kanwar, 2019) with DeepSpeed’s ZeRO-1 optimization (Rajbhandari et al., 2020), reducing memory consumption and allowing for larger model training. This comprehensive setup optimizes efficiency and performance, providing an effective solution for large-scale deep learning model training.
Our training was conducted on 4 NVIDIA DGX servers, each equipped with 8 A100 GPUs, and spanned a period of 4 weeks. Inference was performed on a single NVIDIA DGX server with eight A100 GPUs, utilizing parallel batch inference. During the pre-training phase, we employed a data packing strategy (Krell et al., 2021) and batch dynamic right padding strategy in the instruction tuning phase (Wolf et al., 2020), while the inference phase incorporated a batch left padding strategy (Wolf et al., 2020).
4.2 Evaluation Results
In this section, we provide a detailed analysis of the evaluation and results for both the English and Chinese benchmarks. We discuss task-specific performances and identify key areas of strengths and weaknesses for each model. The following sections present insights for the English and Chinese benchmarks, each highlighting the differences in model capabilities across a variety of NLP tasks.
Figure 3: Comparison of different models’ performance across tasks in the Golden Touchstone benchmark, illustrating average performance for English and Chinese tasks respectively.
4.2.1 Results of English Benchmark
We analyze the English benchmark results shown in Table.5 across task categories shown in left column of Figure.3 such as Sentiment Analysis, Classification, and Summarization. We highlight specific strengths and weaknesses observed in each task and metric.
GPT-4o consistently performs well in tasks like Sentiment Analysis with a Weighted-F1 score of 0.8084 and an accuracy of 0.8093 on the FPB dataset and Multiple Choice, showcasing robustness in understanding sentiment-based and structured questions. However, it struggles significantly in Relation Extraction achieving a Relation-F1 score of 0.1613 on the FinRE dataset and Entity Extraction with an Entity-F1 score of 0.1800 on the NER dataset, suggesting difficulties with detailed information extraction and handling relationships. FinMA-7B excels in Sentiment Analysis on the FPB dataset, achieving a Weighted-F1 score of 0.9400 and an accuracy of 0.9402 but has difficulties in Classification on the FOMC dataset, with a Weighted-F1 score of 0.3988 and an accuracy of 0.4274 and Summarization on the EDTSUM dataset, with Rouge-1, Rouge-2, and BLEU scores all below 0.16, indicating a lack of versatility in broader NLP tasks despite its sentiment strengths. Qwen-2-7B-Instruct performs well in Sentiment Analysis with a Weighted-F1 score of 0.7965 on the FPB dataset but is weak in Summarization on the EDTSUM dataset, achieving a Rouge-2 score of 0.0433 and Question Answering with an RMACC of 0.0270 on the Finqa dataset. Llama-3-8B-Instruct stands out in Stock Movement Prediction achieving a Weighted-F1 score of 0.5116 but struggles in Entity Extraction achieving an Entity-F1 score of 0.2973, highlighting its capacity in structured tasks but limitations in text understanding. FinGPT-8B-lora shows consistent underperformance, especially in Relation Extraction (FinRE, Relation-F1: 0.0100) and Summarization (EDTSUM, BLEU: 0.0592), indicating significant gaps in training these complex financial tasks, necessitating targeted model improvements. The results of many models on lengdingclub are not good enough, this is because of the specialization of the dataset, if the model has not been trained on a similar dataset, the results from zero-shot inference are definitely not good, which emphasizes the importance of instruction fine-tuning for large models in finance. Touchstone-GPT, built on the strong foundation of Qwen-2-7B, utilizes high-quality financial pre-training and instruction tuning datasets through continuous pre-training and post-training, resulting in competitive performance across most tasks.
In summary, the analysis reveals that while GPT-4o and FinMA-7B perform well in sentiment tasks, Llama-3-8B-Instruct has strengths in stock movement prediction, but it’s still not at a practical level. All models, including Touchstone-GPT, have limitations in handling broader NLP tasks.
4.2.2 Results of Chinese Benchmark
We present an analysis of the Chinese benchmark results across tasks such as Sentiment Analysis, Entity Extraction, and Summarization. Specific task metrics are discussed to highlight strengths and weaknesses.
GPT-4o excels in Question Answering with an RMACC of 0.6578 on the FinQa-CN dataset and Entity Extraction achieving an ORMACC of 0.6867 on the FinESE-CN dataset but struggles in Classification with a score of 0.3303 on the FinNL-CN dataset and Relation Extraction with a score of 0.2754 on the FinRE-CN dataset, showing strength in detailed extraction but limitations in handling complex relationships. Qwen-2-7B-Instruct shows moderate but balanced performance, with strengths in Multiple Choice achieving a Weighted-F1 score of 0.7230 on the FinEval dataset but weaknesses in Relation Extraction achieving an ORMACC of 0.1330 and Entity Extraction achieving an ORMACC of 0.3678. Llama-3-8B-Instruct performs particularly well in Stock Movement Prediction (Weighted-F1: 0.4929) but is weak in Sentiment Analysis with a score of 0.4262 on the FinFe-CN dataset and Classification with a score of 0.0747, indicating a less capability in detailed financial text understanding. CFGPT1-7B-Full struggles in most tasks, particularly in Classification with a score of 0.0894 and Relation Extraction (0.0678), with low scores across Summarization metrics with a BLEU score of 0.0238 on the FinNA-CN dataset, implying a need for enhanced adaptability and specialized training. DISC-FinLLM-Full shows moderate potential in Entity Extraction with 0.4346 ORMACC but lacks in Relation Extraction (0.1182) and Classification (0.0011), suggesting opportunities for targeted optimization to enhance broader capabilities. As mentioned above, Touchstone-GPT, based on the Qwen-2-7B model, leverages carefully curated financial pre-training data and instruction tuning through continuous pre-training and post-training, achieving competitive results across most tasks.
In conclusion, while GPT-4o and Llama-3-8B-Instruct have areas of strength in Entity Extraction and Stock Movement Prediction, the models overall, including Qwen-2-7B-Instruct, CFGPT1-7B-Full, and DISC-FinLLM-Full, exhibit considerable gaps in handling relational and classification tasks in Chinese, highlighting the challenges of processing complex Chinese financial text. The need for targeted model improvements remains evident for these tasks.
4.3 Analysis of Evaluation Results
In this overall analysis, we synthesize insights from both the English and Chinese benchmarks by evaluating the strengths and weaknesses of each model, as well as examining the performance across different types of tasks and datasets.
From the perspective of individual models, GPT-4o shows strong performance in sentiment analysis and structured tasks like multiple choice, indicating robust general language understanding capabilities. However, its weakness lies in relation extraction and detailed entity extraction, which require detailed understanding of complex financial relations. FinMA-7B stands out in sentiment tasks but lacks versatility, especially in question answering and summarization, likely due to the absence of targeted training for diverse NLP challenges. Qwen-2-7B-Instruct has a balanced yet modest performance, doing well in sentiment analysis but struggling significantly in question answering and summarization, which suggests a need for more specialized post-training. Llama-3-8B-Instruct excels in stock movement prediction, but shows limitations in tasks requiring deep linguistic understanding, such as entity and relation analysis. The metrics of FinGPT-8B-lora indicating that the current level of domain-specific tuning is insufficient for complex financial tasks. Finally, DISC-FinLLM-Full and CFGPT1-7B-Full demonstrate moderate strengths in entity extraction tasks but lack the robustness needed for broader NLP capabilities, revealing significant gaps in financial language comprehension.
From a task perspective, we observe that Sentiment Analysis generally yields high scores across most models, particularly for the English benchmark, indicating that sentiment understanding, even in financial contexts, is relatively well addressed by these models. In contrast, Relation Extraction and Question Answer in financial domain exhibit notably lower performance, especially for the Chinese benchmark. These results suggest that capturing financial relationships and classifying detailed financial statements pose greater challenges, requiring more sophisticated training datasets or better model architectures. The LendingClub dataset in Classification is a specialized dataset in the field of risk control, requiring more targeted fine-tuning to achieve good results. Stock Movement Prediction also shows low performance across most models, with only a few models such as Llama-3-8B-Instruct demonstrating relatively moderate performance, but it is still practically unusable, highlighting the inherent difficulty of this task. Market prediction relying solely on news information is likely insufficient; volume-price data and factor analysis can provide more comprehensive information. However, current large language models are unable to process these inputs, which is a significant area of future research. Summarization also stands out as a weak area for most models, with consistently low BLEU and Rouge scores, reflecting the challenges in generating concise, coherent summaries of complex financial text.
Overall, the insights suggest that while models like GPT-4o, FinMA-7B, and Touchstone-GPT have particular strengths in sentiment analysis and some structured tasks, the overall capability to handle comprehensive financial NLP tasks remains limited. Most models require targeted improvements, especially for relation extraction, summarization, question answering and stock movement prediction in both English and Chinese contexts. This calls for more domain-specific training and the development of specialized datasets that focus on capturing the detailed and often complex financial language, which is crucial for advancing the performance of financial large language models. Furthermore, while Touchstone-GPT demonstrates competitive performance across various tasks due to its robust pre-training and instruction tuning, ongoing refinements and specialized tuning efforts are needed to address specific deficiencies observed in tasks such as summarization, relation extraction, question answering and stock movement prediction.
Table 5: Performance metrics of financial large language models across english tasks like Sentiment Analysis, Classification, and Summarization. Models include GPT-4o, Llama-3-8B, Qwen-2-7B, FinMA-7B, FinGPT-8B, and Touchstone-GPT. The best results of each dataset are marked in bold.
Task Dataset Metrics GPT-4o FinMA-7B full Qwen-2-7B Instruct Llama-3-8B Instruct FinGPT-8B lora Touchstone GPT
Sentiment Analysis FPB Weighted-F1 0.8084 0.9400 0.7965 0.7631 0.2727 0.8576
ACC 0.8093 0.9402 0.8000 0.7660 0.3072 0.8557
Fiqa-SA Weighted-F1 0.8106 0.8370 0.6726 0.7515 0.5885 0.8591
ACC 0.7702 0.8340 0.5957 0.7064 0.5872 0.8638
Classification Headlines Weighted-F1 0.7857 0.9739 0.7278 0.7006 0.4516 0.9866
ACC 0.7931 0.9739 0.7252 0.7004 0.4331 0.9866
FOMC Weighted-F1 0.6603 0.3988 0.6112 0.4904 0.2758 0.8788
ACC 0.6794 0.4274 0.6210 0.5625 0.2702 0.8790
lendingclub Weighted-F1 0.6730 0.1477 0.5938 0.5943 0.5480 0.9783
MCC 0.1642 -0.6218 0.1714 0.1670 -0.1120 0.9297
Entity Extraction NER Entity-F1 0.1800 0.6200 0.2875 0.2973 0.0231 0.6993
Relation Extraction FinRE Relation-F1 0.1613 0.0054 0.1083 0.0540 0.0100 0.5331
Multiple Choice CFA Weighted-F1 0.7700 0.2200 0.6697 0.5800 0.3993 0.7497
ACC 0.7700 0.2400 0.6700 0.5800 0.3800 0.7500
Summarization EDTSUM Rouge-1 0.1675 0.1566 0.1466 0.1467 0.0622 0.5254
Rouge-2 0.0556 0.0491 0.0433 0.0429 0.0085 0.3446
Rouge-L 0.1069 0.1060 0.0857 0.0930 0.0412 0.4705
BLEU 0.1192 0.1361 0.0999 0.1085 0.0592 0.4512
Question Answering Finqa RMACC 0.1037 0.0497 0.0270 0.0470 0.0110 0.2258
Convfinqa RMACC 0.2540 0.0953 0.0644 0.1477 0.0772 0.5053
Stock Movement Prediction DJIA Weighted-F1 0.4241 0.3211 0.2744 0.5116 0.2171 0.4396
ACC 0.4648 0.3291 0.4372 0.5101 0.2211 0.4749
Table 6: Performance metrics of financial large language models across chinese tasks like Sentiment Analysis, Classification, and Summarization. Models include GPT-4o, Llama-3-8B, Qwen-2-7B, CFGPT-7B, DISC-FinLLM, and Touchstone-GPT. The best results of each dataset are marked in bold.
Task Dataset Metrics GPT-4o Qwen-2-7B Instruct Llama-3-8B Instruct CFGPT1-7B Full DISC-FinLLM Full Touchstone GPT
Sentiment Analysis FinFe-CN Weighted-F1 0.6593 0.6274 0.3633 0.2528 0.4177 0.7888
ACC 0.6500 0.6436 0.4891 0.2732 0.4292 0.7936
Classification FinNL-CN ORMACC 0.3303 0.0622 0.0747 0.0894 0.0011 0.8360
Entity Extraction FinESE-CN ORMACC 0.6867 0.3678 0.3088 0.3863 0.4346 0.9074
Relation Extraction FinRE-CN RMACC 0.2754 0.1330 0.1296 0.0678 0.1182 0.6541
Multiple Choice FinEval Weighted-F1 0.7364 0.7230 0.4432 0.3543 0.4288 0.7361
ACC 0.7353 0.7235 0.4471 0.3529 0.4294 0.7353
CPA Weighted-F1 0.6312 0.6957 0.3421 0.3543 0.3451 0.9238
ACC 0.6309 0.6960 0.3504 0.3553 0.3518 0.9238
Summarization FinNA-CN Rouge-1 0.3197 0.3326 0.3477 0.1018 0.3486 0.5526
Rouge-2 0.1434 0.1597 0.1702 0.0263 0.1678 0.3603
Rouge-L 0.2511 0.2644 0.2802 0.0650 0.2997 0.5214
BLEU 0.1423 0.1541 0.1672 0.0238 0.1885 0.3944
Question Answering FinQa-CN RMACC 0.6578 0.5043 0.4540 0.1126 0.3949 0.9214
FinCQa-CN RMACC 0.4765 0.3422 0.3787 0.2714 0.2134 0.8552
Stock Movement Prediction AStock Weighted-F1 0.5007 0.4906 0.4903 0.4631 0.4142 0.4003
ACC 0.5017 0.4915 0.4956 0.4888 0.4144 0.5587
5 Conclusion and Future Work
In this study, we introduce the Golden Touchstone benchmark, the inaugural structured and comprehensive bilingual benchmark specifically designed for English-Chinese financial NLP. This benchmark encompasses a wide array of financial NLP tasks, including Natural Language Understanding (NLU) and Natural Language Generation (NLG) across eight categories: Sentiment Analysis, Classification, Entity Extraction, Summarization, Stock Market Prediction, Question Answering, Relation Extraction, and Multiple Choice. By leveraging existing high-quality open-source financial datasets, we curated representative datasets and selected appropriate evaluation metrics for each task category. Utilizing these resources, we conducted extensive evaluations of current models such as GPT-4o and prominent open-source financial LLMs, including FinGPT and FinMA, thereby establishing performance benchmarks for financial LLMs within bilingual contexts. Moreover, we contributed to the community by open-sourcing Touchstone-GPT, a robust financial LLM that employs a two-stage training approach and has demonstrated superior input-based inference capabilities on the Golden Touchstone benchmark compared to GPT-4o. Our open-source initiative provides a bilingual English-Chinese evaluation framework aimed at fostering the sustainable development of LLMs in a multilingual financial environment.
Despite these advancements, the benchmark currently exhibits certain limitations, including a limited range of NLG tasks and a focus solely on single-modality. Future enhancements will include the integration of additional NLG tasks, such as extended text generation for financial report analysis and more sophisticated sentiment assessments. Furthermore, we plan to expand the benchmark to cover other financial sectors such as insurance, cryptocurrency, and futures trading, thus broadening the scope and applicability of financial LLM assessments across diverse scenarios. Also, the performance of Touchstone-GPT on specific tasks within the Golden Touchstone benchmark, particularly in stock market prediction, requires further improvement. Our subsequent research will explore the incorporation of agent-based and retrieval-augmented generation (RAG) methods to augment the model’s capabilities in numerical computation and real-time news analysis. Additionally, we aim to venture into multimodal modeling, integrating visual data and time-series data for tasks such as financial time-series forecasting, financial chart analysis, and content generation.
Appendix A Examples of instruction construction
This Table.7 provides a detailed overview of the instruction construction examples for various financial NLP tasks, categorized by task type and language. Each example includes the instruction, input, and output components, demonstrating the specific requirements and expected responses for each task. For a more comprehensive understanding of the prompt settings and additional details, please refer to our dataset’s open-source repository on Hugging Face: Golden-Touchstone Dataset.
Table 7: Examples of Instruction Construction for Various Financial Language Tasks, Categorized by Task Type and Language
Task Type Language Instruction Input Output
Sentiment Analysis English What is the sentiment of the following financial post: Positive, Negative, or Neutral? RT @tomhend777 \$MU needs to hold here -Broken for now. Needs big flush. Still not technically oversold so now big bounce yet neutral
Chinese 以下是股民论坛中的一则股民评论,其中包含有感性的情感输出和理性的涨跌预测等内容…… 判断的非常准确,几次T的相当稳妥! 1
Classification English Review the sentence from a central bank’s communiqué…… In their discussion of prices, participants indicated that data over the intermeeting period…… neutral
Chinese 把接下来输入的金融新闻分类为一个或多个与其描述内容相关的类别…… 加拿大皇家银行:将Affirm Holdings(AFRM.O)目标价从175美元下调至127美元。 外国 公司
Entity Recognition English In the sentences extracted from financial agreements in U.S. SEC filings…… There is a default in any agreement to which Borrower or any Guarantor is a party with a third party or parties…… Borrower, PER
Chinese 给定一段文本T,和文本所属的事件类型S,从文本T中抽取指定事件类型S的事件主体…… 文本: 天龙新材关联担保事项未及时披露被监管佳士科技(300193)股东减持900万股 套现近2亿 事件类型: 信批违规 天龙新材
Relation Extraction English What is the relationship between Ivan Glasenberg and Glencore in the context of the input sentence…… The persistent oversupply is "damaging the credibility of the industry," Glencore CEO Ivan Glasenberg said in May. owner_of
Chinese 给定句子和其中的头尾实体,要求你预测头尾实体之间的关系…… 头实体: ISIS 尾实体: 美军 句子: 美军已对ISIS发动<N>次空袭外资石油巨头欲撤离 unknown
Task Type Language Instruction Input Output
Stock Movement Prediction English Please predict the next rise or fall of DJIA Adj based on the next input of the day’s 25 most popular news items…… Top1:WikiLeaks demands answers after Google hands staff emails to US government…… 1
Chinese 在考量了公司的相关公告之后,请根据新闻对股票数据的影响对该公司股票的表现进行分类…… 公司董事长车成聚承诺自本公告日起未来六个月拟增持价值0.5-1.0亿公司股份…… 0
Multiple Choice English Given a text T, and several options, according to the question posed in the text T…… The inventory/sales ratio is most likely to be rising…… C
Chinese 给定一段文本T,和四个选项ABCD,根据文本T中提出的问题从四个选项中选择合适的多个选项作为答案…… 下列选项中责任中心判断一项成本是否可控的条件有( )…… A,B,D
Summarization English You are given a text that consists of multiple sentences…… PORTLAND, Ore., Feb. 17, 2021 /PRNewswire/ – Allied Market Research published a report, titled,"Matcha Tea Market By Product Type…… Matcha Tea Market to Reach \$4.48 Bn, Globally, by 2027 at 7.1%……
Chinese 请对根据接下来的输入的中文短新闻进行摘要总结,请直接开始总结,不需要输出任何解释 美港电讯APP 13日讯,法航荷航集团(Air France-KLM)已开始与波音(BA.N)和空客就可能成为该集团有史以来最大的飞机订单进行谈判…… 波音空客将竞争法航荷航集团史上最大订单
Question Answering English Please answer the given financial question based on the context…… on november 18 , 2014 , the company entered into a collateralized reinsurance agreement with kilimanjaro…… The answer is:0.26685
Chinese 请根据下面提出的一个问题,问题后的材料内会有相应的答案…… 江苏金沙地理信息股份有限公司公司上市事件对应的证券代码是什么?挖贝网10月9日,全国中小企业股转系统公告显示…… 873361
Appendix B Inference Template of Large Language Models
The Table.8 showcases how inference templates vary across different models. It is crucial to select the appropriate template for constructing correct inputs when inferring on the test sets of datasets. An incorrect template can significantly impair the performance of a model. We have observed that the underperformance of some large financial language models in some benchmarks is precisely due to not selecting the appropriate templates for evaluation. For more details of training and inference template, please refer to our open-source code repository on Github: Golden-Touchstone Code.
Table 8: Comparison of Inference Templates Across Different Models for Dataset Evaluation
Model Template
[…]
Llama-3 "<|start_header_id|>system<|end_header_id|>"
[…]
"<|start_header_id|>user<|end_header_id|>"
[…]
FinGPT "Instruction:{{instruction}}"
[…]
DISC-FinLLM "<reserved_102> {{instruction}}{{input}}<reserved_103>"
[…]
Appendix C Typical Case Study Analysis of Typical Financial NLP Tasks
This appendix provides a detailed case study analysis for some typical financial NLP tasks: financial sentiment analysis, text classification, entity extraction, and stock movement prediction. Each analysis is presented in a separate table, categorizing data sets, instructions, inputs, labels, and predictions from multiple models.
Table 9: Detailed Case Study Analysis of Financial Sentiment Analysis on the FiQA-SA dataset, Financial Text Classification on the LendingClub dataset, Financial Entity Extraction on NER dataset, Stock Movement Prediction on DJIA dataset.
Dataset Instruction Input Label Model Prediction
FiQA-SA What is the sentiment of the following financial post: Positive, Negative, or Neutral? @Joerazorback This time I think \$AAPL will miss again. Generally speaking market is near top when everyone thinks that market goes only up. Negative GPT-4o The sentiment of this financial post is Negative. The user expresses doubt about Apple’s performance, suggesting that they think the stock will miss expectations. Additionally, the phrase about the market being "near top" implies a bearish outlook, indicating that the market may decline after its current uptrend.
Qwen-2 The sentiment of the given financial post is Negative. The user is predicting that Apple’s stock price (\$AAPL) will miss or fail to meet expectations, and they also suggest that the market might be nearing its peak when everyone is optimistic about continuous upward movement.
Llama-3 The sentiment of this post is Negative. The author is predicting that Apple’s stock (AAPL) will miss expectations, which is a negative view. Additionally, the author is suggesting that the market is near a top, implying that a downturn is likely, which is also a negative sentiment.
FinGPT Negative
FinMA Negative
Touchstone GPT Negative
LendingClub Assess the client’s loan status based on the following loan records from Lending Club. Respond with only ’good’ or ’bad’. The client has attributes as follows: … The state of Employment Length is 2 years. The state of Home Ownership is RENT. The state of Verification Status is Verified. The state of Annual Income is 83000.0. Good GPT-4o Good
Qwen-2 Good
Llama-3 Good
FinGPT Bad
FinMA Bad
Touchstone GPT Good
NER In the sentences extracted from financial agreements in U.S. SEC filings, identify the named entities. WHEREAS: (A) The Borrower intends to develop a plant in France for the production of solar grade silicon. Borrower, PER; France, LOC GPT-4o France, LOC
Qwen-2 Borrower, PER; France, LOC; solar grade silicon, ORG
Llama-3 Borrower, PER; France, LOC
FinGPT The Borrower intends to develop a plant in France for the production of solar grade silicon
FinMA Borrower, PER; France, LOC
Touchstone GPT Borrower, PER; France, LOC
DJIA Predict the next movement of DJIA Adj based on the day’s top 25 news items. Top1: b’Haitians are so desperate for food…The Japanese comic version of Das Kapital… 0 GPT-4o [Decrease 0]
Qwen2 [Decrease 0]
Llama-3 [Decrease 0]
FinGPT ""
FinMA Rise 1
Touchstone GPT 1
Financial Sentiment Analysis As demonstrated in Table LABEL:tab:fin_sent_analysis, financial sentiment classification is one of the simpler tasks for benchmarks in financial NLP, resulting in high performance across all models tested. General-purpose models (GPT-4o, Qwen-2, Llama-3) provide not only the answer but also a detailed analysis, despite not being specifically fine-tuned on the FiQA-SA dataset. In contrast, specialized models (FinGPT, FinMA, Touchstone-GPT) that have undergone instruction tuning deliver straightforward, direct responses, illustrating their efficiency and focus in domain-specific applications.
Credit Rating Analysis In the classification task using the LendingClub dataset, which poses a challenging credit rating task, the models face a complex array of professional financial information evident in the input fields. Consequently, most models do not perform optimally. Among general models, GPT-4o exhibits the best performance, demonstrating the capabilities of large-scale models. In the realm of specialized financial language models, Touchstone-GPT, with its high-quality instruction tuning, significantly outperforms FinMA and FinGPT, which are only minimally tuned with Lora.
Financial NER In this information extraction task, most models demonstrated an understanding of the task intent and adhered to the instructions, signifying that even in the era of large language models, models like Qwen-2 and Llama-3 actually outperformed GPT-4o. In particular, specialized models such as FinMA and Touchstone-GPT, with more comprehensive instruction tuning, responded accurately and succinctly, highlighting their enhanced capability and focus on domain-specific tasks.
Stock Movement Prediction The Stock Movement Prediction task is one of the most challenging tasks, as it requires models to predict the daily fluctuations of the DJIA based solely on the top 25 news items from social media. From the results in Table.3, it is evident that Llama-3 performed the best, yet it still falls short of practical utility. Even Touchstone-GPT, despite specialized instruction tuning, performed poorly. Our analysis suggests that the sentiment of news items may not reliably predict stock movements and that incorporating quantitative data is essential for achieving practical model performance. Similar conclusions were drawn from experiments with traditional machine learning methods like XGBoost. Nevertheless, aside from simpler tasks like financial sentiment analysis, we also pose challenging tasks such as stock prediction, which are closer to real-world applications, leaving more room for benchmark challenges and exploration. Multimodal fusion of news and quantitative data represents a promising future direction, and we look forward to seeing models excel in these tasks.
Due to the similar performance of models across corresponding task types on the Chinese benchmark, we will not reiterate these comparisons and analysis here.